OpenStudy (anonymous):

Find the equation of the line through the intersection of the lines 3x+2y−8=0,5x−11y+1=0 and parallel to the line 6x+13y=25 L1 = 3x + 2y -8 = 0, L2 = 5x -11y +1 = 0, L3= ? , L4 = 6x + 13y -25= 0

OpenStudy (cwrw238):

first solve the equations of L1 and L2 simultaneously to find the point of intersection

OpenStudy (anonymous):

L1=3x+2y−8=0 L2=5x−11y+1=0 L3=? L4=6x+13y−25=0 (3x+2y-8) + k(5x-11y+1)= 0 ------(i) 3x+2y-8+5kx-11ky+k =0 Arrange and take common: (3+5k)x + (2 +11k) y - 8 +k =0 The slope from this equation is : -(3+5k)/(2-11k) Since L3 is parallel to L4 therefore: Slope of L3 = Slope Of L4 -(3+5k)/(2-11k) = -6/13 39+15k = 12 -66k 65k+66k = 12-39 131k = -27 k= -27/131 Put the value of k in equation (i) (3x+2y-8) + (-27/131)(5x-11y+1)= 0 393x+262y-1048 -135x+297y-27= 0 258x+559y-1075 =0 Divide by 43: 6x +13y-25 =0 ------(ANSWER) ThankYOU :)

OpenStudy (cwrw238):

yes thats correct

OpenStudy (cwrw238):

its easier to solve the 2 equations first to find point of intersection though then use the general form of the straight line y - y1 = m(x -x1)

OpenStudy (anonymous):

I tried to solve it that way but the answer didn't match.

OpenStudy (cwrw238):

Oh! mine did did you get point of intersection to be (2,1)?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

so m = -6/13 , x1 = 2 and y1 = 1 y - 1 = (-6/13)( x - 2) 13y - 13 = -6x + 12 6x + 13y - 25 = 0